Please help and show all work!7. Ibuprofen is the active ingredient of the over-the-counter painkiller Advil. Its structure is shown below. Like other non-10. Palmitic acid, C15H31COOH, is the first fatty acid created in the biochemical process where excess carbohydrates are conv
Answer #1

Initial conc. of weak acid HA ibuprofen = 0.55 M The reaction of HA in water is Let x be the change in the conc. of each spec

Number of moles of Lactic acid = M*V = 2 M * 500 mL = 1000 mmol Number of moles of KOH = 1M* 500 mL = 500 mmol Net moles of a

concentration of plamitic acid (C15H3COOH) = 0.050 M Ką of C15H3COOH = 1.66 x 10-5 Let x be the change in the conc. of each s